OpenStudy (anonymous):

there were 490 altogether in 2 groups. group A consited of only boys and group b constied of only girls. ther were 2 and a half times as many girls as boys. some girls joined group b and foe every 4 boys in group a 32 more boys jioned the group. the total number og girls was then one third the total number of boys. express the number of girls who jioned group b as a fraction of boys who jioned group a

OpenStudy (horsegirl27):

I would suggest you just write all the information down and it will make more sense. Like, write what you know. Then, write an equation to help you solve. I would help you with this, but I gtg. Just try writing it out yourself more simply, that often helps. The solving will be easy.

OpenStudy (anonymous):

i've tried but its too confusing think if someone exlpained it to me i would grasp the idea better

OpenStudy (anonymous):

a = 1/8 A and b = 1/2 * 1/8 a = a/16 and there are 108 boys at the end

OpenStudy (anonymous):

could you further explain it some more please?

OpenStudy (anonymous):

to begin, A = 1/8 A a boys joined A 32/4 = 8, so 8A boys joined the group.so, A = 1/8 A b girls joined B So, now there are B+b = 1/3(9A) = 3A 5/2 A + b = 3A b = 1/2 A b = 1/2 * 1/8 a = a/16 example: A=12 B=30 A/4 = 3, so 3*32 = 96 boys joined A, making 180 boys in all B+b = 108/3 = 36 so, 6 girls joined, or 96/16

OpenStudy (anonymous):

so how is the number of girls who joined group b expressed as a fraction of boys who joined group a

OpenStudy (anonymous):

12/16